OpenStudy (unklerhaukus):

\[=\sum\limits_{r=1}^\infty\frac{(-1)^r}{1-2r}\] \[=1-\frac13+\frac15-\frac17+\dots\]

OpenStudy (unklerhaukus):

\[\large=\frac\pi4\]

OpenStudy (unklerhaukus):

@experimentX

OpenStudy (experimentx):

http://tinyurl.com/d5x9bk9

OpenStudy (experimentx):

\[ \int_0^1{1 \over 1+x^2}dx = \arctan(1) = {\pi \over 4} \]

OpenStudy (experimentx):

\[ f(u) = \int_0^u {1\over 1+x^2}dx = \int_0^u(1-x^2 + x^4 - x^5 + ...)dx \\\] put 'u=1'

OpenStudy (unklerhaukus):

hmm

OpenStudy (anonymous):